# Towards Federated Graph Learning in One-shot Communication

### v6

@@ Abstract
Federated Graph Learning (FGL) has emerged as a promising paradigm for breaking data silos among distributed private graphs. In practical scenarios involving heterogeneous distributed graph data, personalized Federated Graph Learning (pFGL) aims to enhance model utility by training personalized models tailored to client needs. However, existing pFGL methods often require numerous communication rounds under heterogeneous graphs, leading to significant communication overhead and security concerns. While One-shot Federated Learning (OFL) enables collaboration in a single round, existing OFL methods are designed for image-centric tasks and are ineffective for graph data, leaving a critical gap in the field. Additionally, personalized models derived from existing methods suffer from bias, failing to effectively generalize to the minority. To address these challenges, we propose the first One-shot personalized Federated Graph Learning method (O-pFGL) for node classification, compatible with Secure Aggregation protocols for privacy preservation. Specifically, for effective graph learning in one communication round, our method estimates and aggregates class-wise feature distribution statistics to construct a global pseudo-graph on the server, facilitating the training of a global graph model. To mitigate bias, we introduce a two-stage personalized training approach that adaptively balances local personal information and global insights from the pseudo-graph, improving both personalization and generalization. Extensive experiments on 12 multi-scale graph datasets demonstrate that our method significantly outperforms state-of-the-art baselines across various settings.
1 Introduction
@@
Graphs are widely employed to model complex relationships between entities across a variety of domains [20, 25, 3]. While various algorithms and models [16, 26, 50] have been proposed, most of them assume that graph data from different sources are managed centrally. However, collecting and managing such graph data is often costly, impractical, and poses significant privacy risks in many sensitive scenarios [14].
@@
To address the tension between the need for vast datasets and the growing demand for privacy protection, Federated Graph Learning (FGL) has emerged as a promising solution, which enables collaborative training on distributed graph data without requiring centralized management [37, 17]. While FGL shows its potential, it faces two major limitations: (1) Typical FGL struggles to effectively personalize models for joint improvement across heterogeneous clients’ graph data [2]. This heterogeneity arises from variations of both node and structural properties, with adjacent nodes influencing each other, further complicating personalization. (2) FGL requires iterative communication between clients and the server, raising efficiency concerns due to communication overhead [28, 8] and security concerns (e.g. man-in-the-middle attacks [47, 36, 67]). These limitations can be mitigated by integrating personalized federated graph learning (pFGL) and one-shot federated learning (OFL), which offer novel solutions to improve communication efficiency and model personalization.
@@
To address these challenges, we explore the integration of pFGL and OFL. The pFGL methods allow for personalized model training by focusing on individual client graph data. OFL methods limit communication to a single round, thereby reducing the communication overhead and circumventing security risks such as man-in-the-middle attacks [18, 62]. However, both existing pFGL and OFL methods face challenges in FGL. Specifically, they exhibit the following limitations:
@@
L1: Ineffectiveness for graph learning in one communication round. Existing pFGL methods [2, 33, 69] rely on iterative communication for optimization, making them ineffective within one communication round. Most methods also fail to support heterogeneous models, as they primarily rely on aggregating model parameters. Meanwhile, existing OFL methods are incompatible with graph data since they focus primarily on image-centric tasks. The ensemble-based OFL methods [62, 9] can generate images from ensemble models, but struggle with graph data due to the inter-dependencies between nodes. The distillation-based OFL faces the problems of distilling inter-related graphs. The generative-based OFL methods [21, 55, 54] rely on pre-trained generative models (e.g., Stable Diffusion [38]), which are impractical for graph generation.
@@
L2: Incompatibility with Secure Aggregation. Secure Aggregation is a widely used technique [4, 41] designed to safeguard client privacy during the federated aggregation process. However, most pFGL and OFL methods require clients to upload raw model parameters or auxiliary information independently, which complicates the aggregation process (e.g., ensemble or similarity calculation) and hinders the implementation of privacy-preserving weighted averages.
@@
L3: Imbalance between personalization and local generalization. Local graph data often exhibit a significant imbalance in both quantity and topology, as illustrated in Figure 1. Personalized models tend to be biased toward major classes (with larger quantities and high homophily to form compact communities), while neglecting minor classes with fewer samples and often surrounded by nodes from major classes. This imbalance leads to a degradation in model performance, particularly in underrepresented classes. A detailed empirical analysis is provided in Figure 1, which shows that existing personalized methods result in heavy bias toward majority classes (class 3 and class 5, with a total of 180 samples) while failing to adequately represent minority classes (class 1 and class 2, with only 64 samples), highlighting the limitations of current methods in handling imbalanced graph data.
@@
To overcome these limitations, we propose a One-shot personalized Federated Graph Learning method (O-pFGL). Our method supports model heterogeneity and Secure Aggregation protocols for privacy preservation. Specifically, for L1 and L2, clients upload estimated feature distribution statistics, augmented by our proposed homophily-guided reliable node expansion strategy, instead of raw model parameters. The server securely aggregates these statistics to generate a global pseudo-graph, which is subsequently utilized to train personalized models on clients, eliminating the need for direct model parameter exchanges and preventing model theft. To address L3, we propose a two-stage personalized training approach. In the first stage, clients train a global generalized model based on the pseudo-graph, which serves as both a fixed teacher capturing global knowledge and a start-point for subsequent fine-tuning. In the second stage, clients fine-tune the global model on local graph data with the node-adaptive distillation, which leverages the global knowledge from the teacher model to counteract biases toward the majority, ensuring effective learning for the minority and achieving a robust balance between personalization and generalization.
@@
Our method significantly outperforms state-of-the-art baselines through comprehensive experiments across various settings, including different graph scales, partition strategies, learning paradigms, and model architectures. Experimental results demonstrate the superiority of our method in achieving a one-shot personalized federated graph learning solution that balances efficiency, privacy, and personalization. To sum up, our contributions are listed as follows:
@@
Bridge the Gap. To the best of our knowledge, we are the first to investigate the one-shot personalized federated graph learning problem under both graph data heterogeneity and model heterogeneity.
@@
Novel Solution. We propose a novel method that aggregates client statistics rather than model parameters to generate a global pseudo-graph, which is used for two-stage personalized training with node-adaptive distillation. Our method is compatible with Secure Aggregation and protects models’ intellectual property.
@@
SOTA Performance. We conduct extensive experiments on 12 multi-scale graph datasets across various settings and demonstrate the superiority of our method.
2 Related Works
2.1 Federated Graph Learning
@@
From the graph level [52, 43], each client possesses multiple independent graphs (e.g., molecular graphs) and targets at graph-level tasks. From the subgraph level, each client possesses a subgraph of an implicit global graph. FedPUB [2] and FedGTA [33] personalize models by similarity of functional embedding and mixed moments of neighbor features. AdaFGL [32] studies the structure non-IID problem. FedTAD [69] performs data-free distillation on the server. To complete missing connections between subgraphs, FedSage [63] and FedDEP [64] additionally train a neighborhood generator. To better represent the minority in local graph data, FedSpray [15] learns class-wise structure proxies to mitigate biased neighborhoods. But it needs numerous communication rounds for optimization. FedLoG [30] synthesizes global data from clients’ synthetic nodes and generalizes the local training. Detailed description and critical analysis are in the Appendix A.6.
2.2 One-shot Federated Learning
@@
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into: (1) Ensemble-based methods [18, 62, 9], (2) Distillation-based methods [66, 42], (3) Generative-based methods [21, 54, 55]. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning. Detailed description and critical analysis are in the Appendix A.6.
3 Preliminaries
3.1 Notations and Definitions
@@
Consider a federated graph learning setting with clients, where -th client possesses the local undirected graph with edges and nodes. Node features matrix is , where is the feature dimension and is the feature of node . The adjacency matrix is and the augmented normalized adjacency matrix is . The labels of nodes in its local graph are denoted by , where is the number of distinct classes. In the semi-supervised node classification task [31], nodes on the -th client are partitioned into labeled nodes (i.e., nodes in the training set) and unlabeled nodes (i.e., their labels are unknown during training). The corresponding labels are and , respectively. The neighborhood of node in is defined as , where edge connects node and .
Definition 3.1.
Node Degree. For node in an undirected graph , the node degree is the number of edges connected to . Formally, .
Definition 3.2.
Node Homophily. For a labeled node in an undirected graph , the node homophily quantifies the proportion of its labeled neighbors that share the same label:
Definition 3.3.
Accumulated Class Homophily. The accumulated class homophily for class is defined as:
@@
where is the set of labeled nodes belonging to class . Note that considers both the quantity and topology of nodes. A higher indicates nodes of class are numerous and exhibit strong homophily, making them more predictable for a vanilla graph model [31].
Definition 3.4.
Class-aware Distillation Factor.
which can be organized into a vector where the -th entry corresponds to for class . is used to determine node-adaptive distillation factor , as described in Section 4.4.
3.2 Problem Formulation
@@
Assume a situation where a whole graph is distributed on clients, each of which possesses its local non-overlapped subgraph. We formulate the one-shot personalized federated graph learning for the node classification problem as:
@@
4 Proposed Method
4.1 Overall Pipeline
The overall pipeline of our method is illustrated in Figure 2, which comprises 5 steps. (1) Each client estimates the class-wise feature distribution and computes corresponding statistics on its local graph data; (2) Each client uploads its locally estimated class-wise statistics to the server; (3) The server aggregates the uploaded statistics by weighted average to recover the global class-wise feature distribution. Based on the recovered distribution, the server generates a small-size global pseudo-graph; (4) The server distributes the generated global pseudo-graph to clients; (5) Each client performs two-stage personalized training using both local graph data and the downloaded global pseudo-graph, deriving a model with improved personalization and generalization.
Specifically, in the first step, we propose a homophily-guided reliable node expansion strategy to augment the distribution estimation. In the third step, we adapt techniques from graph condensation to facilitate the global pseudo-graph generation. In the fifth step, we address the critical bias issue in personalization by proposing a two-stage personalized training process with node-adaptive distillation. The details are explained in the following sections.
4.2 Feature Distribution Estimation
@@
where is the propagation depth and represents the concatenation. Using the propagated features, we unbiasedly estimate the sample mean and variance of features for labeled nodes in each class:
@@
However, the estimation may lack accuracy due to the limited number of labeled nodes. To address this, we propose a Homophily-guided Reliable node Expansion (HRE) strategy as a plugin to further augment the estimation process on homophilic graphs. In HRE, we first apply Label Propagation [26] to obtain the soft labels for unlabeled nodes. We denote , then identify reliable soft labels based on the following criteria: (C1) High Node Degree: ; (C2) High Confidence: ; (C3) High Class Homophily: , indicating the node is likely to share the same label with its neighbors. A node with more neighbors could reduce the variance in predictions. With these criteria, the reliable nodes set with inferred true labels is:
where is the set of classes corresponding to the top largest values of accumulated class homophily . and are client-defined hyper-parameters. The reliable nodes , along with their inferred labels, can be used to expand the original labeled nodes set . The estimation of feature distribution could be augmented as:
where represents the subset of reliable nodes in belonging to class . This augmentation ensures a more accurate and robust estimation of class-wise feature distributions by incorporating reliable nodes into the computation.
4.3 Global Pseudo-Graph Generation
@@
which are used to guide the global pseudo-graph generation.
Inspired by the graph condensation [35, 51], we generate a small-size global pseudo-graph by aligning its feature distribution with the aggregated global distribution. This pseudo-graph serves as a representative of the overall graph data across all clients. The generation process is illustrated in the top of Figure 2. We first pre-set the number of nodes of each class in the pseudo-graph and initialize the trainable node features from Gaussian noise. The adjacency matrix is derived using a trainable link predictor :
where is Sigmoid function, is a hyper-parameter to control sparsity, and represents concatenation. We implement as a simple MLP. We also discuss other designs in the Appendix A.3.9.
Following [51], to optimize and , we first propagate node features as in Eq. 5. Then we calculate the class-wise sample mean and sample variance of the propagated node features following Eq. 6. Then we calculate the alignment loss:
where is the proportion of nodes belonging to class relative to the total number of nodes. To ensure the smoothness of the pseudo-graph, smoothness loss is applied:
We could optimize and to generate the global pseudo-graph by the overall optimization objective: . Once the optimization finishes, the server distributes to clients for further personalized training.
Privacy-preservation with Secure Aggregation. We demonstrate compatibility with Secure Aggregation protocols in uploading and aggregating . Secure Aggregation is a crucial technique in federated learning that enables the server to compute the sum of large, user-held data vectors securely, without accessing individual client contributions. This property aligns well with the aggregation process in Eq. 9, as it can be expressed as a series of weighted average operations. The calculations of and are inherently weighted average operations. The calculation of , while more complex, can also be decomposed into weighted average operations:
Breaking this down further, the calculation involves the weighted averages of and . Thus, the entire aggregation process for can be conducted without revealing individual client data contributions, making it fully compatible with Secure Aggregation.
4.4 Personalized Training with Node Adaptive Distillation
With the downloaded global pseudo-graph, each client trains its personalized model locally. However, the imbalanced local graph data often results in biased models that perform well on major classes but neglect minor classes. To achieve better personalization (accuracy) and generalization (F1-macro), we propose a two-stage personalized training approach with node-adaptive distillation to leverage both global information in the global pseudo-graph and personal information in the local graph.
@@ 4.4 Personalized Training with Node Adaptive Distillation
Each client trains a generalized model using only the global pseudo-graph, ensuring to capture global, balanced knowledge. The training objective is to minimize:
A copy of is detached as a fixed teacher model for the next stage. denotes the cross-entropy loss. Note that if the server knows the model architecture, an alternative is to conduct this stage on the server. However, we can adapt to more rigorous scenarios where models’ intellectual property needs to be protected. In those cases, it’s necessary to conduct this stage on each client.
@@
where is the cross-entropy loss for supervised fine-tuning on local training data, and is the weighted sum of point-wise KL divergence loss to distill global knowledge from into :
where is a node-specific weighting factor to control the balance between local fine-tuning and global knowledge distillation. To determine for each node, we re-use the soft labels obtained in Section 4.2 and in Eq. 3:
where is a scaling hyper-parameter. Intuitively explained, if a node is predicted to class with a small (e.g. node in Figure 7), it likely belongs to major classes with high homophily. In this case, is small, and primarily learns through supervised fine-tuning; If a node is predicted to class with a large or has a blended prediction (e.g. node in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus should be large to incorporate the corresponding global knowledge from .
@@
| O-pFGL (Ours) | 76.43 1.24 | 61.58 2.16 | 71.61 0.40 | 58.24 0.96 | 82.71 0.03 | 66.40 0.34 | 70.93 0.59 | 36.66 0.22 | 86.63 0.02 | 27.21 0.15 |
5 Experiments
5.1 Experimental Setup
@@
We conduct experiments on 12 datasets, including both transductive learning datasets: Cora, CiteSeer, PubMed [56], ogbn-arxiv, ogbn-products [23], Computers, Photo, CS, Physics datasets [40], and inductive learning datasets: Flickr [59], Reddit [19], and Reddit2 [59]. Table 4 summarizes these datasets. We employ two advanced graph partitions to simulate distributed subgraphs: Louvain-based Label Imbalance Split and Metis-based Label Imbalance Split [34].
@@
We compare our method with 9 state-of-the-art methods, including 5 pFGL methods111We do not include FedSage [63] and AdaFGL [32] since they need extra communication rounds to train neighborhood generators or federated optimization. They cannot train models within one-shot communication.: (1) FedPUB [2], (2) FedGTA [33], (3) FedTAD [69], (4) FedSpray [15], (5) FedLoG [30], and 2 OFL methods: (6) DENSE [62], (7) Co-Boost [9]. We also include (8) Standalone, which locally trains models, and (9) FedAvg with fine-tuning. Detailed descriptions are in the Appendix A.7.
@@
We limit the communication round to a single round. We employ the two-layer GCN [31] with the hidden dimension set to 64. We employ both accuracy(%) and F1-macro(%) metrics. F1-macro averages the F1 scores across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. We report the mean and standard deviation across three runs. Detailed configurations are described in the Appendix A.2.
5.2 Experimental Results
5.2.1 Performance Comparison

### v7

# Personalized One-shot Federated Graph Learning for Heterogeneous Clients

@@ Abstract
Federated Graph Learning (FGL) has emerged as a promising paradigm for breaking data silos among distributed private graphs. In practical scenarios involving heterogeneous distributed graph data, personalized Federated Graph Learning (pFGL) aims to enhance model utility by training personalized models tailored to client needs. However, existing pFGL methods often require numerous communication rounds under heterogeneous graphs, leading to significant communication overhead and security concerns. While One-shot Federated Learning (OFL) enables collaboration in a single round, existing OFL methods are designed for image-centric tasks and are ineffective for graph data, leaving a critical gap in the field. Additionally, personalized models derived from existing methods suffer from bias, failing to effectively generalize to the minority. To address these challenges, we propose the first One-shot personalized Federated Graph Learning method (O-pFGL) for node classification, compatible with Secure Aggregation protocols for privacy preservation. Specifically, for effective graph learning in one communication round, our method estimates and aggregates class-wise feature distribution statistics to construct a global surrogate graph on the server, facilitating the training of a global graph model. To mitigate bias, we introduce a two-stage personalized training approach that adaptively balances local personal information and global insights from the surrogate graph, improving both personalization and generalization. Extensive experiments on 14 diverse real-world graph datasets demonstrate that our method significantly outperforms state-of-the-art baselines across various settings.
Introduction
@@
Graphs are widely employed to model complex relationships between entities across a variety of domains (Hyun, Lee, and Suh 2023; Bang et al. 2023). While various algorithms and models (Gasteiger, Bojchevski, and Günnemann 2018; Wu et al. 2020) have been proposed, most of them assume that graph data from different sources are managed centrally. However, collecting and managing such graph data is often costly, impractical, and poses significant privacy risks in many sensitive scenarios (Fu et al. 2022).
@@
To address the tension between the need for vast datasets and the growing demand for privacy protection, Federated Graph Learning (FGL) has emerged as a promising solution, which enables collaborative training on distributed graph data without requiring centralized management (Pan, Zhu, and Chu 2023; Gu et al. 2023). While FGL shows its potential, it faces two major limitations: (1) Typical FGL struggles to effectively personalize models for joint improvement across heterogeneous clients’ graph data (Baek et al. 2023). This heterogeneity arises from variations of both node and structural properties, with adjacent nodes influencing each other, further complicating personalization. (2) FGL requires iterative communication between clients and the server, raising efficiency concerns due to communication overhead (Kairouz et al. 2021; Dai et al. 2022) and security concerns (e.g., man-in-the-middle attacks (Wang et al. 2020)). These limitations can be mitigated by integrating personalized federated graph learning (pFGL) and one-shot federated learning (OFL), which offer novel solutions to improve communication efficiency and model personalization.
@@
To address these challenges, we explore the integration of pFGL and OFL. The pFGL methods allow for personalized model training by focusing on individual client graph data. OFL methods limit communication to a single round, thereby reducing the communication overhead and circumventing security risks such as man-in-the-middle attacks (Guha, Talwalkar, and Smith 2019; Zhang et al. 2022). However, both existing pFGL and OFL methods face the following challenges and limitations:
@@
L1: Ineffectiveness for graph learning in one communication round under heterogeneity. Existing pFGL methods (Baek et al. 2023; Li et al. 2024b) rely on iterative communication, making them ineffective within one communication round, especially under heterogeneous data. A recent work, GHOST (Qian et al. 2025), explores training a generalized model in one-shot communication. But it’s not robust under severe non-IIDness and does not address personalization issues, as confirmed in our experiments. Furthermore, such parameter-aggregation-based methods inherently fail to support heterogeneous client models. Meanwhile, existing OFL methods are incompatible with graph data, as they primarily focus on image-centric tasks. Ensemble-based OFL methods (Zhang et al. 2022; Dai et al. 2024) can generate images from ensemble models but struggle with graph data due to the inter-dependency between nodes. Distillation-based OFL faces problems distilling interrelated graphs. Generative-based OFL methods (Heinbaugh, Luz-Ricca, and Shao 2023; Yang et al. 2024b, a) rely on pre-trained generative models (e.g., Stable Diffusion (Rombach et al. 2022)), which are impractical for graph generation.
@@
L2: Incompatibility with Secure Aggregation. Secure Aggregation is a widely used technique (Bonawitz et al. 2017) designed to safeguard client privacy during uploading and aggregation in federated learning. However, most existing methods require clients to upload raw model parameters or auxiliary information independently, which complicates the process (e.g., similarity calculation or model ensemble) and hinders the implementation of privacy-preserving uploading and aggregation. The individual information could be revealed in the communication (Zhu et al. 2024a).
@@
L3: Imbalance between personalization and local generalization. Local graph data often exhibits a significant imbalance in both quantity and topology. Personalized models tend to be biased toward major classes (e.g., with larger quantities and high homophily to form compact communities), while neglecting minor classes with fewer samples and often surrounded by nodes from major classes. This imbalance leads to a degradation in model performance, particularly in underrepresented classes. A detailed empirical analysis is provided in Figure 1, which shows that existing methods result in heavy bias toward majority classes (class 3 and class 5, with a total of 180 samples) while failing to adequately represent minority classes (class 1 and class 2, with only 64 samples), highlighting the limitations of current methods in handling imbalanced graph data.
@@
To overcome these limitations, we propose a One-shot personalized Federated Graph Learning method (O-pFGL). Our method supports model heterogeneity and Secure Aggregation protocols for privacy preservation. Specifically, for L1 and L2, clients upload estimated feature distribution statistics, augmented by our proposed homophily-guided reliable node expansion strategy, instead of raw model parameters. The server securely aggregates these statistics to generate a global surrogate graph, which is subsequently utilized to train personalized models on clients, eliminating the need for direct model parameter exchanges and further protecting the client’s model intellectual property. To address L3, we propose a two-stage personalized training approach. In the first stage, clients train a global generalized model based on the surrogate graph, which serves as both a fixed teacher capturing global knowledge and a starting point for subsequent fine-tuning. In the second stage, clients fine-tune the global model on local graph data with the node-adaptive distillation, which leverages the global knowledge from the teacher model to counteract biases toward the majority, ensuring effective learning for the minority and achieving a robust balance between personalization and generalization.
@@
Our method significantly outperforms state-of-the-art through comprehensive experiments across various settings, including different graph scales, homophilic and heterophilic, learning paradigms, and model architectures. Experimental results demonstrate our method’s superiority in achieving a personalization-generalization balanced model within one-shot communication, considering both data and model privacy. To sum up, our contributions are:
@@
New Problem: One-shot Personalized FGL. We are the first to tackle the problem of personalization in one-shot federated graph learning under dual challenges of data and model heterogeneity, and ensuring both privacy.
@@
Novel Solution: Our solution firstly aggregates client statistics rather than model parameters to generate a global surrogate graph that captures global distribution. Leveraging this, we then introduce two-stage personalized training with node-adaptive distillation to strike an effective balance between personalization and generalization.
@@
SOTA Performance: We conduct extensive experiments on 14 diverse real-world graph datasets and demonstrate that our method consistently and significantly outperforms existing baselines across various settings.
Related Works
Federated Graph Learning
@@
From the graph level (Xie et al. 2021; Tan et al. 2023), each client possesses multiple independent graphs and targets at graph-level tasks. From the subgraph level, each client possesses a subgraph of an implicit global graph. FedPUB (Baek et al. 2023) and FedGTA (Li et al. 2024b) personalize models by similarity of functional embedding and mixed moments of neighbor features. AdaFGL (Li et al. 2024a) studies the structure non-IID problem. FedTAD (Zhu et al. 2024b) performs data-free distillation on the server. FedGM (Zhang et al. 2025) adopts gradient matching to synthesize graphs and aggregate the gradients. GHOST (Qian et al. 2025) trains aligned proxy models, and the server integrates the knowledge from uploaded proxy models. To complete missing connections between subgraphs, FedSage (Zhang et al. 2021) and FedDEP (Zhang et al. 2024b) additionally train a neighborhood generator. FedStruct (Aliakbari, Östman et al. 2024) decouples the structure learning and node representation learning. To better represent the minority in local graph data, FedSpray (Fu et al. 2024) learns class-wise structure proxies to mitigate biased neighborhoods. FedLoG (Kim et al. 2025) synthesizes global data from clients’ synthetic nodes and generalizes the local training. Detailed description and analysis are in the Appendix H.
One-shot Federated Learning
@@
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into: (1) Ensemble-based methods (Guha, Talwalkar, and Smith 2019; Zhang et al. 2022; Dai et al. 2024), (2) Distillation-based methods (Zhou et al. 2020; Song et al. 2023), (3) Generative-based methods (Heinbaugh, Luz-Ricca, and Shao 2023; Yang et al. 2024a, b). However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning. Detailed description and analysis are in the Appendix H.
Preliminaries
Notations and Definitions
@@
Consider a FGL setting with clients, where -th client possesses the local undirected graph with edges and nodes. Node features matrix is , where is the feature dimension and is the feature of node . is the adjacency matrix. The labels of nodes are denoted by , where is the number of distinct classes.
In the semi-supervised node classification task (Kipf and Welling 2016), nodes on the -th client are partitioned into labeled nodes (i.e., training set) and unlabeled nodes (i.e., test set). The corresponding labels are and , respectively. The neighborhood of node in is , where edge connects node and . denotes the degree of node . We define the node homophily for node as:
and the class homophily for class is:
@@
where is the set of labeled nodes belonging to class . considers both the quantity and topology of nodes. A higher indicates nodes of class are numerous and exhibit strong homophily, making them more predictable for a vanilla graph model.
Problem Formulation
@@
We formulate the one-shot personalized federated graph learning for the node classification over clients with non-overlapped distributed subgraphs as:
@@
Proposed Method
The overall pipeline of our proposed method is illustrated in Figure 2. The details are explained in the following sections.
Feature Distribution Estimation
@@
where is propagation depth and represents concatenation. is augmented normalized adjacency matrix. Using propagated features, we estimate the unbiased sample mean and variance of features for labeled nodes in each class:
@@
Our estimation is compatible with both homophilic and heterophilic graphs. For heterophilic graphs, the concatenation in Eq 4 ensures the ego-information and high-order information are not conflated. This aligns with the core principles of MixHop (Abu-El-Haija et al. 2019), of which effectiveness has been widely proven on heterophilic graphs. For homophilic graphs, we further propose a plug-and-play Homophily-guided Reliable node Expansion (HRE) strategy to enhance the accurate estimation when the labeled nodes are scarce. In HRE, we first apply Label Propagation (Iscen et al. 2019) to obtain the soft labels for unlabeled nodes. We denote , then identify reliable soft labels based on the following criteria: (C1) High Node Degree: ; (C2) High Confidence: ; (C3) High Class Homophily: , indicating the node is likely to share the same label with its neighbors. A node with more neighbors could reduce the variance in predictions. With these criteria, the reliable nodes set with inferred true labels is:
where is the set of classes with top largest values of . and are client-defined hyper-parameters. The reliable nodes , along with their inferred labels, can expand the original labeled nodes set . The estimation in Eq. 5 could be augmented as:
where is the subset of reliable nodes in belonging to class . This augmentation ensures a more accurate estimation of class-wise feature distributions by incorporating reliable nodes into the computation.
Global Surrogate Graph Generation
@@
Notably, the recovered global statistics remain unbiased regardless of the degree of data non-IIDness.
With recovered global distribution, we generate a small-sized global surrogate graph by aligning its feature distribution with global distribution. The surrogate graph serves as a representative of the latent overall graph data across all clients. The generation process is illustrated at the top of Figure 2. We first pre-set the number of nodes in the surrogate graph and initialize the learnable node features with Gaussian noise. The adjacency matrix is derived using a learnable link predictor :
where is Sigmoid function, is a hyper-parameter to control sparsity, and is concatenation. We implement as an MLP. We also discuss other designs in the Appendix D.4.
Following (Xiao et al. 2024), to optimize and , we first propagate node features as in Eq. 4. Then we calculate the class-wise sample mean and sample variance of the propagated node features following Eq. 5. Then we calculate the alignment loss:
where is the proportion of nodes belonging to class relative to the total number of nodes. To ensure the smoothness of the surrogate graph, a smoothness loss is applied:
We optimize and by the overall optimization objective: . Once the optimization finishes, the server distributes to clients for further personalized training.
Privacy-preservation with Secure Aggregation. We demonstrate compatibility with Secure Aggregation protocols in uploading and aggregating . Secure Aggregation is a crucial technique in federated learning that enables the server to compute the sum of large, user-held data vectors securely, without accessing individual client contributions. This property aligns with the aggregation process in Eq. 8, as it can be expressed as a series of weighted average operations. The calculations of and are inherently weighted average operations. The calculation of can also be decomposed into weighted average operations:
Breaking this down further, the calculation involves the weighted averages of and . Thus, our entire aggregation process is compatible with off-the-shelf Secure Aggregation protocols and would not reveal individual client data distribution. We detail the implementation with pseudo code in Appendix B.
Personalization with Node Adaptive Distillation
With the downloaded global surrogate graph, each client trains its personalized model locally. However, the imbalanced local graph data often results in biased models that perform well on major classes but neglect minor classes. To achieve better personalization (accuracy) and generalization (F1-macro), we propose a two-stage personalization training approach with node-adaptive distillation, which leverages both global information in the global surrogate graph and personal information in the local graph.
@@
Each client trains a generalized model using the global surrogate graph, ensuring to capture global, balanced knowledge. The training objective is to minimize:
A copy of is detached as a fixed teacher model for the next stage. denotes the cross-entropy loss. Note that if the server knows the model architecture, an alternative is to conduct this stage on the server. However, we can adapt to more rigorous scenarios where models’ intellectual property needs to be protected. In these cases, it’s necessary to conduct this stage on each client.
@@
where is cross-entropy loss for fine-tuning on local data, and is the weighted sum of point-wise KL divergence loss to distill global knowledge from into :
where is a node-specific weighting factor to control the balance between local fine-tuning and global knowledge distillation. To determine for node , we consider both the node property and class property. We first define the class-aware distillation weight vector where:
is defined in Eq 2. Then we use node ’s soft label by Label Propagation and to calculate :
where is a scaling hyper-parameter. Intuitively explained, if a node is predicted to class with a small (e.g., node in Figure 6 in Appendix A), it likely belongs to major classes with high homophily. In this case, is small, and primarily learns through fine-tuning; If a node is predicted to class with a large or has a blended prediction (e.g. node in Figure 6 in Appendix A), it likely belongs to minor classes or has low homophily. These nodes are often neglected during fine-tuning. Thus should be large to incorporate the corresponding global knowledge from .
@@
| O-pFGL | 76.431.24 | 61.582.16 | 71.610.40 | 58.240.96 | 82.710.03 | 66.400.34 | 70.930.59 | 36.660.22 | 86.630.02 | 27.210.15 |
Privacy Discussion
Our method provides privacy for both client models and data. Model Privacy: By not sharing model parameters, our method inherently mitigates model-based attacks (e.g., model inversion and membership inference attacks (Zhu et al. 2024a)). This design crucially allows clients to use their proprietary models, fostering collaboration without exposing intellectual property. Data Privacy: Client data is protected twofold. First, the shared statistics are computed via a non-injective operation (Eq. 7), making it computationally infeasible to reverse-engineer specific data samples. Second, our framework is fully compatible with Secure Aggregation protocols, which prevent the server from accessing any individual client’s statistics.
| Method | Actor | genius | ||
| Acc. | F1 | Acc. | F1 | |
| Standalone | 58.770.28 | 20.640.22 | 96.271.38 | 53.301.71 |
| FedAvg | 60.010.21 | 21.170.12 | 94.791.95 | 53.301.71 |
| FedPUB | 63.430.06 | 21.110.17 | 95.581.55 | 51.61.65 |
| FedGTA | 62.580.07 | 16.050.06 | 92.231.67 | 47.670.26 |
| FedTAD | 60.320.12 | 21.300.11 | OOM | |
| FedSpray | 65.810.46 | 20.130.18 | 95.341.71 | 53.330.36 |
| FedLoG | 61.490.82 | 20.530.38 | OOM | |
| FedGM | 58.990.09 | 20.940.19 | 96.880.27 | 53.920.28 |
| GHOST | 58.940.09 | 20.700.07 | OOM | |
| DENSE | 59.880.11 | 21.050.02 | 96.810.13 | 53.970.10 |
| Co-Boost | 60.020.53 | 21.110.22 | 96.730.11 | 53.800.12 |
| O-pFGL | 67.540.32 | 22.630.26 | 97.050.17 | 54.220.17 |
Experiments
Experimental Setup
@@
We conduct experiments on 14 datasets, including Cora, CiteSeer, PubMed (Yang, Cohen, and Salakhudinov 2016), ogbn-arxiv, ogbn-products (Hu et al. 2020), Computers, Photo, CS, Physics datasets (Shchur et al. 2018), Flickr (Zeng et al. 2019), Reddit (Hamilton, Ying, and Leskovec 2017), Reddit2 (Zeng et al. 2019), Actor, and genius (Luan et al. 2024). Table 3 summarizes these datasets. For Actor and genius datasets, we employ Label non-IID partition (Qian et al. 2025) with . For the others, we employ Louvain/Metis-based Label Imbalance Split (Li et al. 2024c) to simulate distributed subgraphs.
@@
Our baselines include 7 FGL methods111We do not include FedSage (Zhang et al. 2021), AdaFGL (Li et al. 2024a), and FedStruct (Aliakbari, Östman et al. 2024) since they need extra communication rounds before federated optimization. They cannot train models within one-shot communication.: (1) FedPUB (Baek et al. 2023), (2) FedGTA (Li et al. 2024b), (3) FedTAD (Zhu et al. 2024b), (4) FedSpray (Fu et al. 2024), (5) FedLoG (Kim et al. 2025), (6) FedGM (Zhang et al. 2025), (7) GHOST (Qian et al. 2025) and 2 OFL methods: (8) DENSE (Zhang et al. 2022), (9) Co-Boost (Dai et al. 2024). We also include (10) Standalone, which locally trains models, and (11) FedAvg with fine-tuning. Detailed descriptions of these baselines are in the Appendix I.
@@
We conduct experiments over a single communication round. We employ the two-layer GCN with a hidden dimension of 64. We employ both accuracy(%) and F1-macro(%) metrics. F1-macro provides a more robust measure of the model’s generalization under imbalanced data. We report the mean and standard deviation across three runs. In our O-pFGL, the global surrogate graph is set small, with its size not exceeding 300 nodes across all datasets. Detailed configurations are in the Appendix C.
Experimental Results
Performance Comparison